OpenStudy (mathmath333):

remainder for 3^124/17

OpenStudy (ikram002p):

3^3 = 10 mod 17 ok ?

OpenStudy (mathmath333):

book says remainder 4

OpenStudy (ikram002p):

its 4 :P but do u have any method to solve it ?

OpenStudy (mathmath333):

(ax+1)^n/a gives remainder 1

OpenStudy (mathmath333):

(81)^31/17

OpenStudy (mathmath333):

(85-4)^31

ganeshie8 (ganeshie8):

nice :)

ganeshie8 (ganeshie8):

but that wont be useful cuz the last term would be (-4)^31

ganeshie8 (ganeshie8):

try ikram's method atleast you will be able to take powers of 10 easily :)

OpenStudy (mathmath333):

(3^7)^17 will leave the same remainder when divided by 17 as 3^7, so we can address 3^7 · 3^5 = 3^12 And since 3^12 has remainder 4 when divided by 17, so will 3^124 PLEASE EXPLANE THE ABOVE

ganeshie8 (ganeshie8):

id like to try ikram's method first : \[\large 3^{124} = 3^{3\times 40 + 4} = 3^4*27^{40} \equiv 3^4*10^{40} \]

OpenStudy (mathmath333):

THE original question was 54^124/17 which reduces to 3^124/17

ganeshie8 (ganeshie8):

\[\large \begin{array} \\ 3^{124} = 3^{3\times 40 + 4} = 3^4*27^{40} &\equiv 3^4*10^{40} \\ &\equiv 3^4*100^{20}\\ &\equiv 3^4*(-2)^{20} \\ &\equiv 3^4*(-2)^{4\times 5} \\ &\equiv 3^4*(16)^{ 5} \\ &\equiv 3^4*(-1)^{ 5} \\ &\equiv -3^4 \\ \end{array}\]

ganeshie8 (ganeshie8):

see if that looks okay so far ^^

OpenStudy (mathmath333):

can the (-4)^31/ 17 solved further ?

ganeshie8 (ganeshie8):

Oh yes ! that gives you answer much faster !

ganeshie8 (ganeshie8):

\[\large (-4)^{31} = (-4)^{2\times 15 + 1} = -4 (-4)^{2\times 15} = -4(16)^{15}\]

ganeshie8 (ganeshie8):

\[\large \equiv -4(-1)^{15}\]

OpenStudy (mathmath333):

thats easy

ganeshie8 (ganeshie8):

\[\large \equiv -4(-1) = 4\] yep :)

OpenStudy (ikram002p):

nico

OpenStudy (ikram002p):

i just rememberd that 3^16=1 mod 17 xD

OpenStudy (ikram002p):

but thought he only use like calculation stuff

ganeshie8 (ganeshie8):

i think only ppl who take number theory knw about fermat

OpenStudy (ikram002p):

xD but its cool isnt it :P i mean whats the point of calculation while we have fermat <3

ganeshie8 (ganeshie8):

3^16 = 1 mod 17 may throw off other students, @mathmath333 do you want to know why 3^16/17 gives you 1 as remainder ?

OpenStudy (ikram002p):

=o

ganeshie8 (ganeshie8):

exactly ! if he knows fermat already, there is no point

OpenStudy (mathmath333):

well ,i dont know the modulo method lol

ganeshie8 (ganeshie8):

Fair enough ! binomial expansion is powerful, but it works only in special cases... modulo method is more universal

OpenStudy (mathmath333):

can u give link for learning modulo method

ganeshie8 (ganeshie8):

if u have time, i can teach u... it only takes 30 minutes to see the beauty of congruences !

OpenStudy (ikram002p):

:o lol that would be cool to see how u do it in 30 mnts

OpenStudy (mathmath333):

ya

ganeshie8 (ganeshie8):

ok start the clock

ganeshie8 (ganeshie8):

it all starts with \(1\) definition and \(5\) theorems : Definition : \(\large a \equiv b \mod n\) means \(\large a-b\) is divisible by \(\large n\)

OpenStudy (mathmath333):

whats that 3 lines means

OpenStudy (ikram002p):

new post :o

ganeshie8 (ganeshie8):

for example : \[\large 11\equiv 1 \mod 10\] cuz \(\large 11-1\) is divisible by \(\large 10\)

OpenStudy (mathmath333):



ganeshie8 (ganeshie8):

you can read \(\large \equiv \) as `congruent to`

OpenStudy (anonymous):

congruency i guess

ganeshie8 (ganeshie8):

yes ^^

OpenStudy (ikram002p):

time pasing :P continue

OpenStudy (mathmath333):

ok

ganeshie8 (ganeshie8):

let me know once you're happy with the definition

OpenStudy (mathmath333):

ya i got it defn

ganeshie8 (ganeshie8):

fre more examples : \[\large 7 \equiv 1 \mod 6\] \[\large 100 \equiv 1 \mod 99\] \[\large 23 \equiv -1 \mod 24\]

OpenStudy (mathmath333):

ok

ganeshie8 (ganeshie8):

all above examples must make sense before we dive into the \(5\) theorems

ganeshie8 (ganeshie8):

basically, we only need to stick to above definition to understand ANYTHING in congruences/modulo world

ganeshie8 (ganeshie8):

ready for first theorem ? :)

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

Theorem 1 : if \(\large a\equiv b \mod n\) and \(\large c\equiv d \mod n\), then : \(\large a+c \equiv b+d \mod n\) \(\large ac \equiv bd \mod n\)

OpenStudy (mathmath333):

ok

ganeshie8 (ganeshie8):

Is the statement of theorem clear ?

ganeshie8 (ganeshie8):

lets see an example

OpenStudy (mathmath333):

yaa

ganeshie8 (ganeshie8):

do you want to try and cook up an example ? :)

OpenStudy (mathmath333):

yes of cos

ganeshie8 (ganeshie8):

good, make some example..

ganeshie8 (ganeshie8):

I'll give you first example, maybe you can create the next example :)

ganeshie8 (ganeshie8):

we know that : \(3 \equiv 1 \mod 2\) \(2 \equiv 0 \mod 2\) so, by Theorem 1 : \(3+2\equiv 1+0 \mod 2\)

OpenStudy (mathmath333):

7=1 mod 6 and 13=7 mod 6

ganeshie8 (ganeshie8):

yes, thats a good example ^^ apply Theorem 1, what do u get ?

OpenStudy (mathmath333):

7+13=6+7

ganeshie8 (ganeshie8):

dont forget the mod symbol : 7+13=1+7 mod 6

OpenStudy (mathmath333):

okk

ganeshie8 (ganeshie8):

Theorem1 just says this : Sum of two numbers leaves the same reamainder as the sum of remainders of individual numbers

ganeshie8 (ganeshie8):

7+13=1+7 mod 6 20 = 8 mod 6 is this statement correct ? verify the definition

OpenStudy (mathmath333):

yaa

OpenStudy (mathmath333):

wait

ganeshie8 (ganeshie8):

20 = 8 mod 6 means 20-8 is divisible by 6 thats the definition, right ?

ganeshie8 (ganeshie8):

Is 20-8 divisible by 6 ?

OpenStudy (mathmath333):

nooooooo

ganeshie8 (ganeshie8):

why not

ganeshie8 (ganeshie8):

20-8 = ?

OpenStudy (mathmath333):

it leaves remainder

ganeshie8 (ganeshie8):

20-8 = 12 are you saying 12 leaves a remainder when divided by 6 ?

OpenStudy (mathmath333):

oh sry

OpenStudy (mathmath333):

i tjought 20-6

ganeshie8 (ganeshie8):

Theorem 1 is a theorem ! and a theorem can NEVER fail.

ganeshie8 (ganeshie8):

lets do one more example using theorem 1

OpenStudy (mathmath333):

yaa

ganeshie8 (ganeshie8):

5 = 1 mod 4 16 = 0 mod 4 apply Theorem 1